Circular No. 7464

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 2000cv IN PGC 39222 S. Jha, P. Challis, and R. Kirshner, Harvard-Smithsonian Center for

Astrophysics, report: “A spectrum (range 550–750 nm) of SN 2000cv, obtained by P. Berlind on July 24.15 UT with the F. L. Whipple Observatory 1.5-m telescope (+ FAST spectrograph), shows that it is likely a type-Ia supernova near maximum light. Though the wavelength range is limited, the spectrum shows broad absorption features at 581 and 627 nm that we identify as Si II. Cross-correlation of the host-galaxy spectrum with an emission-line template yields a recession velocity of 6120 ± 20 km/s. The expansion velocity of the supernova, measured with the Si II (rest 635.5 nm) profile, is 10 200 km/s.”

COMET C/1999 S4 (LINEAR) C. Lisse and D. Christian, Space Telescope Science Institute; K. Den-

nerl, Max-Planck-Institut für Extraterrestrische Physik; F. Marshall, R. Mushotzky, R. Petre, and S. Snowden, NASA/Goddard Space Flight Center; H. Weaver, Johns Hopkins University; B. Stroozas, University of California; and S. Wolk, Harvard-Smithsonian Center for Astrophysics, report the first detection of x-ray line emission due to charge exchange between cometary neutrals and solar-wind minor ions using Chandra and EUVE: “Using a 960-s ACIS-S observation of comet C/1999 S4 on July 14.20 UT, the comet was detected with a rate in the S3 chip of 0.3 count/s, with a total integral flux of 8 × 10^{−13} erg s^{−1} cm^{−2} over 0.2–0.7 keV and a total x-ray luminosity of 6 × 10^{14} erg/s. The ACIS-S spectrum showed a strong line at 570 eV detected at greater than 10σ, due to charge exchange to O VII. Other lines due to charge exchange are also present (e.g., N VI, N VII at 300–500 eV, O VIII at 650 eV) at lower S/N. The best fit to the preliminary spectrum is the MEKAL multiple-line emission model (Mewe et al. 1986, A.Ap. Suppl. 65, 511) using solar elemental abundances with enhanced oxygen and nitrogen abundances, and a plasma temperature of 0.17 keV. The EUVE Lexan B count rate at July 14.21 was 0.06 count/s in 5400 s, for an equivalent luminosity of 1 × 10^{15} erg/s at 0.16 keV. The observed emission was found to be highly time variable on the order of hours, enhanced by a strong solar flare propagating radially from the sun.”

Visual m_{1} estimates: July 20.91 UT, 7.0 (N. Biver, Ablis, France, 7×50 binoculars); 21.86, 6.7 (D. V. Fedotov, Kharkov, Ukraine, 7×50 binoculars); 23.87, 6.2 (J. Kysely, Vlasim, Czech Republic, 7×50 binoculars); 24.92, 6.2 (A. Baransky, Kiev, Ukraine, 12×80 binoculars).

2000 July 25  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2000 CBAT  Daniel W. E. Green